OpenStudy (anonymous):

If cot theta = 1.5 and theta is in quadrant 3, what is the value of sin theta?

OpenStudy (aravindg):

Use the right angled triangle.

OpenStudy (aravindg):



OpenStudy (aravindg):

Once you get the hypotenuse, find sin theta.

OpenStudy (campbell_st):

well start by recognising that cot = 1/tan so if cot = 3/2 then tan =2/3 use pythagoras to find the hypotenuse... and you know the opposite side is 2... then you can find sin.... remember 3rd quadrant sin is negative